A red flowered plant is crossed with a white flowered plant. All of the F1 plants produce red flowers. When the F1 plants are allowed to self-pollinate, 389 of the F2 plants have red flowers and 141 have white flowers. What are the genotypes of the parental and F1 generation plants?

A. P:Red = RR, white = rr; F1 = Red RR and Rr

B. P:Red =Rr, white = rr; F1 = Red Rr and rr

C. P: Red = RR, white = rr; F1 = Red Rr only

D. P: Red = Rr, white = rr; F1 = Red Rr only
Answer #1

Option A

Parents = RR (red) × rr (white) = F1 = Rr (red)

Selfing of F1 = Rr × Rr

Gametes R r
R RR (red) Rr (red)
r Rr (red) rr (white)

F2 = RR, Rr = red, rr = white

Please rate.